OpenStudy (anonymous):

[5.02] 2x + 1 5 or -3x > -9 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution

OpenStudy (ash2326):

@pbking101 what's the inequality sign here 2x+1 ___________ 5

OpenStudy (anonymous):

greater then or equal to

OpenStudy (ash2326):

Cool, so we have \[2x+1 \ge 5\] or \[-3x>-9\] Let's solve it first, then I'll plot this on graph. Did you try to solve this?

OpenStudy (anonymous):

i dont know how to can u help and all u have to do is describe how u would plot the grapth u dont have to grapth it

OpenStudy (ash2326):

Let's solve the inequalities first \[2x+1\ge 5\] Subtract both sides by 1 \[2x \ge 4\] Divide both sides by 2 \[x \ge 2\] Do you understand this part?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

can u plzz solve it for me i would really appreciate it this is my last question

OpenStudy (ash2326):

You didn't get my solution ?

OpenStudy (anonymous):

no can u plzz just right the answer out full i only have 5 more min to turn the test in please

OpenStudy (ash2326):

@pbking101 I can't help you, if you are taking a test right now. Please refer to the code of conduct http://openstudy.com/code-of-conduct Please try to do the test yourself :)

OpenStudy (anonymous):

to late now its over canu u just show me how i should have done it